6. If 10P30-0.8 and 20Ps0-0.6, find the probability that of three persons now aged 30, at least two will die between ages 40 and 50.] 0.104
Answer #1

Let T be future lifetime random variable, then

10p30 = P(T > 40) = 0.8

20p30 = P(T > 50) = 0.6

Probability that the person at age 30 will die between ages 40 and 50 = P(T > 40) - P(T > 50) = 0.8 - 0.6 = 0.2

Number of people (X) die out of 3 persons will follow binomial distribution with parameters n = 3, p = 0.2

By Binomial distribution, probability that at lest two will die between ages 40 and 50 is,

P(X \ge 2) = P(X = 2) + P(X = 3)

= 3C2 * 0.22 * (1 - 0.2)1 +  3C3 * 0.23 * (1 - 0.2)0

= 3 * 0.04 * 0.8 + 0.008

= 0.104